Jack is x years old. His brother Dan is 5 years younger. In 3 years time the sum of their ages will be 15. Write an equation in terms of x using the sum of their ages in 3 years time.
Answer by rm29924(97)

Jack's current age = x
Dan's current age = x - 5
Jack's age in 3 years = x + 3
Dan's age in 3 years = x - 5 + 3
The sum of their ages in 3 years will be 15.
The equation for the sum of their ages in 3 years is : x + 3 + x - 5 + 3 = 15
if you need to solve the equation combine like terms
2x + 1 = 15
subtract 1 from both sides
2x = 14
divide 2 from both sides to get x by itself
x = 7
Jack's age in 3 years = 7 + 3 = 10
Dan's age in 3 years = 7 - 5 + 3 = 5